OpenStudy (anonymous):

Given that Begin equation . . . begin fraction . . . 2 over 3 . . . end fraction . . . times a equals 16 . . . end equation, which of the following illustrates how the multiplicative inverse property can be used to find the value of a?

OpenStudy (anonymous):

2/3a = 16 multiplication inverse property (3/2) 2/3a = (3/2)16 a = 48/2 a = 24 check.. 2/3(24) = 16 48/3 = 16 16 = 16 (correct)